Ambulance

You are walking along a small country road one foggy morning and come to an intersection. While you are crossing, you hear an ambulance siren. As a paramedic, you knowthat the siren sounds at a frequency of 7.07 kHz. But, as an amateur musician with perfect pitch, you can tell that the frequency you are hearing is actually 7.87 kHz.Is the ambulance moving toward or away from you? What speed (MPH) is the ambulance traveling?
Answer #1
First observation is, the ambulance is going away from you. You hear the frequency too low.
Assuming 343m/sec. as sound speed as you have not specified,:
343 - (343 x (7.07/7.87)) = 20.05m/sec. ambulance speed, away from you.
answered by: Williams
Answer #2
7.87/7.07 = 343/(343-x)

x = 34.86
answered by: Matt